OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 – x2 – 6x ?

OpenStudy (accessdenied):

Set f(x) = 0 x^3 - x^2 - 6x = 0; solve for x's

OpenStudy (bahrom7893):

x(x^2-x-6)=0

OpenStudy (anonymous):

i dont understand

OpenStudy (bahrom7893):

pull out an x: x(x^2-x-6)=0 This means that either x=0, or: x^2-x-6=0

OpenStudy (anonymous):

thanks

OpenStudy (bahrom7893):

just solve x^2-x-6=0 to get the rest of x values.